# ReMoS: 3D Motion-Conditioned Reaction Synthesis for Two-Person Interactions

### v2

@@ 5.5 Qualitative Results and Applications
Fig. 1 shows qualitative results of ReMoS on ReMoCap and highlights the synthesized hand interactions. We note that our actor and reactor are interchangeable depending on the scenario. That is, depending on the character driving the interaction at a given time, we can swap the actor and the reactor to produce relevant reactive motions. Fig. 5 (top) shows a visual comparison of ReMoS with the baselines for one frame of Lindy Hop motion. ReMoS synthesizes reaction with the most plausible alignment with the actor’s motion.
Inspired by the applicability of diffusion-based methods for motion editing and controlling, we demonstrate how ReMoS can also be used as an interactive motion editing tool to control the reactor’s motion in a given manner. In Fig. 5 (bottom) we show results from two different motion editing applications, namely, pose completion with controlled joints and motion in-betweening. We discuss more details in the appendix.
@@ 6 Conclusion
Acknowledgements.
This research was supported by the EU Horizon 2020 grant Carousel+ (101017779), and the ERC Consolidator Grant 4DRepLy (770784).
Appendix A ReMoCap Dataset Analysis
Our proposed ReMoCap dataset covers two types of motion, namely the Lindy Hop dance and the martial art technique of Ninjutsu (see Sec. 4 of the main paper).
@@ Appendix A ReMoCap Dataset Analysis
The Lindy Hop part of the dataset consists of $8$ dance sequences captured at $50$ fps, each around $7.5$ minutes long, resulting in around $174.2\mathrm{K}$ motion frames. We had $4$ trained dancers, $2$ males (identified as A and B) and $2$ females (identified as C and D), participate in the Lindy Hop motion capture. We pair the dancers as (A, C), (B, D), (A, D), and (B, C). Of these pairings, (A, D) contains dance sequences not performed by the other three pairs (in terms of twists and maneuvers). We also capture multiview RGB videos at $50$ fps from $116$ camera views for each sequence, which can benefit two-person pose reconstruction work in the future. We show samples from these videos in Fig. 6. From these sequences, almost $145.2\mathrm{K}$ frames have a hand-in-hand contact between the two dancers with a contact threshold of $50$ mm between the finger joints of the two dancers. By increasing the contact threshold to $100$ mm, the number of frames where the two dancers have contact increases to $147.5\mathrm{K}$. We also analyze the pose diversity of the Lindy Hop dance motion by checking how many distinct poses are obtained, as done by Guo et al. [24]. The percentage of distinct poses among all the poses in Lindy Hop motion data is around $28\%$ when the threshold for diversity is set to $100$ mm.
@@
The Ninjutsu part of the dataset consists of $79$ sequences each captured at $25$ fps. The sequences vary in length with a total number of around $99.8\mathrm{K}$ motion frames resulting in around $66.5$ minutes of motion. We had $5$ trained, male Ninjutsu artists participate in the Ninjutsu motion capture. We pair them in all possible combinations and ask them to perform different variations of motion. Along with the 3D pose, we also capture multiview RGB videos at $25$ fps using $116$ cameras. We show samples from these videos in Fig. 7. From these sequences, almost $81\mathrm{K}$ frames have contact-based interactions between the two performers, where the closest distance between any joints is $50$ mm. The pose diversity of the Ninjutsu motion data is around $18\%$ when the threshold for diversity is set to $100$ mm.
Figure 6: Samples from the Lindy Hop motion capture for the ReMoCap dataset. We show multi-view RGB samples with corresponding 3D poses from our Lindy Hop motion capture performed by trained dancers. Lindy-hop requires coordination between the two dancers, while also allowing individual dancers the freedom to perform their own motions. This makes it suitable for testing our reactive motion synthesis approach.
Figure 7: Samples from the Ninjutsu motion capture for the ReMoCap dataset. We show multi-view RGB samples with corresponding 3D poses from the Ninjutsu motion capture performed by trained artists. In contrast to existing martial arts datasets [54, 81], we include finger joint MoCaps and moves of varying interaction complexity.
Appendix B Dataset and Baseline Preparation
@@ B.1 Dataset Preparation
We split the dataset such that motions captured from dancer pairs (A, C), (B, D), and (B, C) are in our training set, and motions captured from pair (A, D) are in our test set. We downsample each motion sequence to $20$ fps and filter the frames where the dancing partners have hand-to-hand contact between the actors’ and reactors’ finger joints. We represent each character using $27$ body joints and $22$ hand joints. We convert the 3D joint angle representations into joint positions using forward kinematics and then convert them to root local representations, as explained in Sec. 5.1 of the main paper. For training, we use a sequence length of $20$ frames.
@@
We divide the whole dataset into roughly $3:1$ train-test ratio and take $28$ sequences of diverse attacking and maneuvering motions for testing, and the rest for training. We downsample each motion sequence to $10$ fps, and filter out the frames where the pairs are more than $1$ meter of each other. We represent each character using $27$ body joints and $22$ hand joints. We convert the 3D joint angle representations into joint positions using forward kinematics and then convert them to root local representations, as explained in Sec. 5.1 of the main paper. For training, we use a sequence length of $50$ frames.
@@
The ExPI dataset consists of $2$ pairs of professionals performing acrobatics and Lindy Hop aerial sequences. It consists of $16$ different acrobatic actions. Each couple consists of a leader and a follower. We aim to synthesize the motions of the followers as they react to the leaders’ movements. We use the common action split proposed by the original authors [24], and split the dataset into train and test sets such that all the actions performed by (A, B) are in the train set and all the actions performed by (C, D) are in the test set. We represent each subject using $16$ joints (omitting the ‘lhead’ and ‘rhead’ joints) and convert the global 3D joint positions given in the dataset to root relative joint representations, as explained in Sec. 5.1 of the main paper. Since the ExPI dataset does not have hand motions, we only train with body motions and forego the hand diffusion stage. We train ReMoS for about $20\mathrm{K}$ iterations on the ExPI dataset using the Adam optimizer [36], with a base learning rate of $10^{-5}$ and a batch size of $32$.
@@
The 2C dataset consists of full-body motions of kickboxing actions performed by pairs of participants. The interactions include motions such as kicking and punching, with diverse reactions such as avoiding and being hit. We use the pose sequence of the leading character, who throws the punches and kicks, as the acting sequence for our model. We aim to synthesize the full body motion of the reacting character, who is blocking or avoiding the moves, as our output. Following the split of MixNMatch [20], we use a roughly $3:1$ train-test ratio to train our method. Each character contains $25$ joints and we convert the 3D joint angle representations into joint positions using forward kinematics and then convert them to root relative joint position representations, as explained in Sec. 5.1 of the main paper. Since the 2C dataset does not have hand motions, we only train with body motions and forego the hand diffusion stage. We train ReMoS for about $25\mathrm{K}$ iterations on the 2C dataset using the Adam optimizer [36], with a base learning rate of $10^{-5}$ and a batch size of $16$.
@@
The InterHuman dataset consists of human-human interactions for daily motions, such as passing objects, greeting, and communicating, and professional activities, such as, Taekwondo, Latin dance, and boxing. It consists of a total of $7,779$ motions with $22$ joints per person. We randomly select the pose sequence of one of the characters as the acting sequence for each motion to train our model. We aim to synthesize the full body motion of the corresponding other character in each motion as our output. We follow the split of InterGen [42] for our experiments. Since the InterHuman dataset does not have hand motions, we only train with body motions and forego the hand diffusion stage. We train ReMoS for about $45\mathrm{K}$ iterations on the InterHuman dataset using the Adam optimizer [36], with a base learning rate of $10^{-5}$ and a batch size of $64$.
B.2 Baseline Preparation
As we mention in Sec. 5.2 of the main paper, we use InterFormer [11], MixNMatch [20], ComMDM [53], RAIG [64] and InterGen [42] as baselines. We describe how we use each of these methods in our setting.
@@ B.2 Baseline Preparation
InterFormer consists of a transformer network with temporal and spatial attentions. It takes an input acting sequence $Y$ and encodes it with spatial and temporal self-attention. It also needs the initial pose of the reactor $X$ and predicts the subsequent frames of the reactor in an autoregressive manner. It uses information from skeletal adjacency matrices and an interaction distance module that provides information on the interactions. We use the normalization technique mentioned in Sec. 5.1 of the main paper to normalize the actor’s and the reactor’s body poses. We train InterFormer on an NVIDIA RTX A4000 GPU for about $20\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $10^{-5}$ and a batch size of $128$. We use $207$ dimensional latent embedding and $6$ layers in the transformer decoder with $3$ heads to calculate the attention.
@@
MixNMatch proposes an end-to-end framework to synthesize stylized reactive motion informed by multi-hot action labels. For interaction mixing, it generates a reaction combining different classes of reactive styles according to the multi-label indicator. For interaction matching, it generates the reactive motion corresponding to the interaction type and the input motion. Our setting is similar to interaction matching, where we input the acting sequence into the model and synthesize the reactive sequence. We mask out the action label defining the interaction type from the input and train the reactor’s motion $X$ based on the actor’s motion $Y$. We use the normalization technique mentioned in Sec. 5.1 of the main paper to normalize the actor’s and the reactor’s body poses. We train MixNMatch on an NVIDIA RTX A4000 GPU for about $3.6\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $10^{-5}$ and a batch size of $16$. We use $256$ LSTM neurons for each spatial slice and $1200$ for the attention layer.
@@
Role-Aware Interaction Generation (RAIG) is a diffusion-based model that learns two-person interactions by generating single-person motions for a designated role. The role is supplied in the form of textual annotations, which are translated into active and passive voice to ensure the text is consistent with each role. The model generates interactions with two transformers that share parameters, and a cross-attention module connecting them. The active and passive voice descriptions are input to the corresponding transformers responsible for generating the actor and the reactor. The transformers consist of cross-attention modules both for language and motion. To use RAIG as a baseline for our annotation-free setting, we mask out the cross-attention module for the language in both the transformers and train to generate two-person motions unconditionally. We normalize the interactions as described in the original paper [64]. We train RAIG on an NVIDIA RTX A4000 GPU for about $20\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $2^{-4}$ and a batch size of $32$. We use $512$ dimensional latent embedding and $8$ attention blocks. During inference, we freeze the transformer that learns the actor’s motion $Y$. The other transformer generates the reactor’s motion $X$, being influenced by the actor’s ground truth motion.
@@
InterGen is a diffusion-based approach that generates two-person motions from text prompts. It was originally trained conditioned on rich textual annotations. It uses cooperative denoisers with novel weight-sharing and a mutual attention mechanism to improve interaction between two persons. To use it as a baseline in our annotation-free setting, we mask out the text embeddings from the model input, and train InterGen to generate two-person motions (both actor and reactor) unconditionally. We use the non-canonical motion representation proposed in the original paper [42]. During inference, we use the customization used in InterGen for person-to-person generation. We take a single-person motion (the actor’s motion $Y$) as input, and freeze it during the forward diffusion process. The frozen weights from the first person propagate into the model and the model uses the ground truth actor’s motions to reconstruct the second person’s motion (the reactor’s motion $X$). We train InterGen on an NVIDIA RTX A4000 GPU for about $30\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $10^{-4}$, a cosine LR scheduler, and a batch size of $64$.
Appendix C Additional Results
Table 5: Quantitative evaluation on the InterHuman dataset [42]. We compare ReMoS with state-of-the-art motion synthesis methods on the InterHuman [42] dataset. $\downarrow$: lower is better, $\uparrow$: higher is better, $\rightarrow$: values closer to GT are better. Bold indicates best.

### v3

@@ 5.5 Qualitative Results and Applications
Fig. 1 shows qualitative results of ReMoS on ReMoCap and highlights the synthesized hand interactions. We note that our actor and reactor are interchangeable, i.e., depending on the character driving the interaction at a given time, we can swap the actor and the reactor to produce the relevant reactive motions. Fig. 5(a) shows a visual comparison of ReMoS with the baselines for one frame of Lindy Hop motion. ReMoS synthesizes reaction with the most plausible alignment with the actor’s motion. We provide detailed visual results in our supplementary video. Inspired by the applicability of diffusion-based methods for motion editing and controlling, we demonstrate how ReMoS can also be used as an interactive motion editing tool to control the reactor’s motion as desired. In Figs. 5(b) and 5(c), we show results from two different motion editing applications, namely, pose completion with controlled joints and motion in-betweening. We discuss more details in the appendix.
(a) Visual comparison of ReMoS with GT and baselines.
(b) Pose-completion application.
(c) Motion in-betweening application.
Figure 5: Qualitative Results and Applications. We show some visual results and the application of ReMoS as a motion editing tool. (a) The reactor (in blue) synthesized by ReMoS has the most plausible alignment with the actor (in red) compared to the baselines. (b) We manually control the right-hand wrist joint of the reactor and let ReMoS synthesize the remaining body joints conditioned on the actor. (c) ReMoS synthesizes the reactor’s motion in-between the start and end frames.
@@ 6 Conclusion
Acknowledgements
This research was supported by the EU Horizon 2020 grant Carousel+ (101017779), and the ERC Consolidator Grant 4DRepLy (770784). We thank Marc Jahan, Christopher Ruf, Michael Hiery, Thomas Leimkühler and Sascha Huwer for helping with the Ninjutsu data collection, and Noshaba Cheema for helping with the Lindy Hop data collection. We also thank Janis Sprenger and Duarte David for helping with the motion visualizations.
Appendix 0.A Additional Details of Loss Functions
Kinematic Loss Terms.
We describe the details of the velocity, acceleration, bone length and foot sliding losses loss terms from Eqn. 10 in the main paper. To improve the temporal consistency of the motion [68], we minimize the joint velocities and joint accelerations between two consecutive frames of the ground-truth reactive motions, $X$, and the synthesized reactive motions, $\hat{X}$, defined as
$\mathcal{L}_{vel}$ $=\frac{1}{N-1}\sum_{n=0}^{N-1}\left\|\left(X^{n+1}-X^{n}\right)-\left(\hat{X}^{n+1}-\hat{X}^{n}\right)\right\|^{2}_{2},$ (0.A.1)
$\mathcal{L}_{acc}$ $=\frac{1}{N-2}\sum_{n=0}^{N-2}\left\|\left(X^{n+2}-2X^{n+1}+X^{n}\right)-\left(\hat{X}^{n+2}-2\hat{X}^{n+1}+\hat{X}^{n}\right)\right\|^{2}_{2},$ (0.A.2)
where $N$ is the total number of frames.
Additionally, we introduce a bone length consistency loss, $\mathcal{L}_{bone}$, to ensure that the synthesized reactor joint positions satisfy the skeleton consistency [42]. We define this loss as
$\mathcal{L}_{bone}=\left\|\mathbf{B}\left(X\right)-\mathbf{B}\left(\hat{X}\right)\right\|^{2}_{2},$ (0.A.3)
Further, foot sliding is a common artifact in motion synthesis [56, 57]. We constrain this by ensuring that the toe joint in contact with the ground plane has zero velocity. We use a binary foot contact loss [68, 67] on the foot joints of the synthesized pose to ensure that the output motion does not slide across the ground plane, defined as
$\mathcal{L}_{foot}=\frac{1}{N-1}\sum_{n=0}^{N-1}\left\|\left(\hat{X}^{n+1}-\hat{X}^{n}\right)\cdot\hat{\mathbbm{1}}^{n}_{foot}\right\|^{2}_{2},$ (0.A.4)
where $\hat{\mathbbm{1}}^{n}_{foot}\in\left\{0,1\right\}$ is the foot-ground contact indicator for the synthesized reactive motion $\hat{X}^{n}$ at each frame $n$.
Appendix 0.B ReMoCap Dataset Analysis
Our proposed ReMoCap dataset covers two types of motion, namely the Lindy Hop dance and the martial art technique of Ninjutsu (see Sec. 4 in the main paper).
@@ Appendix 0.B ReMoCap Dataset Analysis
The Lindy Hop part of the dataset consists of $8$ dance sequences captured at $50$ fps, each around $7.5$ minutes long, resulting in around $174.2\mathrm{K}$ motion frames. We had $4$ trained dancers, $2$ males (denoted A and B) and $2$ females (denoted C and D), participate in the Lindy Hop motion capture. We pair the dancers as (A, C), (B, D), (A, D), and (B, C). Of these pairings, (A, D) contains dance sequences not performed by the other three pairs (in terms of twists and maneuvers). We also capture multiview RGB videos at $50$ fps from $116$ camera views for each sequence, which can benefit two-person pose reconstruction work in the future. We show samples from these videos in Fig. 0.B.1. From these sequences, almost $145.2\mathrm{K}$ frames have a hand-in-hand contact between the two dancers with a contact threshold of $50$ mm between the finger joints of the two dancers. By increasing the contact threshold to $100$ mm, the number of frames where the two dancers have contact increases to $147.5\mathrm{K}$.
@@
The Ninjutsu part of the dataset consists of $79$ sequences each captured at $25$ fps. The sequences vary in length with a total number of around $99.8\mathrm{K}$ motion frames resulting in around $66.5$ minutes of motion. We had $5$ trained, male Ninjutsu artists participate in the Ninjutsu motion capture. We pair them in all possible combinations and ask them to perform different variations of motion. Along with the 3D pose, we also capture multiview RGB videos at $25$ fps using $116$ cameras. We show samples from these videos in Fig. 0.B.2. From these sequences, almost $81\mathrm{K}$ frames have contact-based interactions between the two performers, where the closest distance between any joints is $50$ mm.
Figure 0.B.1: Samples from the Lindy Hop motion capture for the ReMoCap dataset. We show multi-view RGB samples with corresponding 3D poses from our Lindy Hop motion capture performed by trained dancers. Lindy-hop requires coordination between the two dancers, while also allowing individual dancers the freedom to perform their own motions. This makes it suitable for testing our reactive motion synthesis approach.
Figure 0.B.2: Samples from the Ninjutsu motion capture for the ReMoCap dataset. We show multi-view RGB samples with corresponding 3D poses from the Ninjutsu motion capture performed by trained artists. In contrast to existing martial arts datasets [55, 83], we include finger joint motion capture and moves of varying interaction complexity.
Appendix 0.C Dataset and Baseline Preparation
@@ 0.C.1 Dataset Preparation
We split the dataset such that motions captured from dancer pairs (A, C), (B, D), and (B, C) are in our training set, and motions captured from pair (A, D) are in our test set. We downsample each motion sequence to $20$ fps and filter the frames where the dancing partners have hand-to-hand contact between the actors’ and reactors’ finger joints. We represent each character using $27$ body joints and $22$ hand joints. We convert the 3D joint angle representations into joint positions using forward kinematics and then convert them to root local representations, as explained in Sec. 5.1 in the main paper. For training, we use a sequence length of $20$ frames.
@@
We divide the whole dataset into roughly $3:1$ train-test ratio and take $28$ sequences of diverse attacking and maneuvering motions for testing, and the rest for training. We downsample each motion sequence to $10$ fps, and filter out the frames where the pairs are more than $1$ meter of each other. We represent each character using $27$ body joints and $22$ hand joints. We convert the 3D joint angle representations into joint positions using forward kinematics and then convert them to root local representations, as explained in Sec. 5.1 in the main paper. For training, we use a sequence length of $50$ frames.
@@
The ExPI dataset consists of $2$ pairs of professionals performing acrobatics and Lindy Hop aerial sequences. It consists of $16$ different acrobatic actions. Each couple consists of a leader and a follower. We aim to synthesize the motions of the followers as they react to the leaders’ movements. We use the common action split proposed by the original authors [24], and split the dataset into train and test sets such that all the actions performed by (A, B) are in the train set and all the actions performed by (C, D) are in the test set. We represent each subject using $16$ joints (omitting the ‘lhead’ and ‘rhead’ joints) and convert the global 3D joint positions given in the dataset to root relative joint representations, as explained in Sec. 5.1 in the main paper. Since the ExPI dataset does not have hand motions, we only train with body motions and forego the hand diffusion stage. We train ReMoS for about $20\mathrm{K}$ iterations on the ExPI dataset using the Adam optimizer [36], with a base learning rate of $10^{-5}$ and a batch size of $32$.
@@
The 2C dataset consists of full-body motions of kickboxing actions performed by pairs of participants. The interactions include motions such as kicking and punching, with diverse reactions such as avoiding and being hit. We use the pose sequence of the leading character, who throws the punches and kicks, as the acting sequence for our model. We aim to synthesize the full body motion of the reacting character, who is blocking or avoiding the moves, as our output. Following the split of MixNMatch [20], we use a roughly $3:1$ train-test ratio to train our method. Each character contains $25$ joints and we convert the 3D joint angle representations into joint positions using forward kinematics and then convert them to root relative joint position representations, as explained in Sec. 5.1 in the main paper. Since the 2C dataset does not have hand motions, we only train with body motions and forego the hand diffusion stage. We train ReMoS for about $25\mathrm{K}$ iterations on the 2C dataset using the Adam optimizer [36], with a base learning rate of $10^{-5}$ and a batch size of $16$.
@@
We report additional results on the InterHuman dataset in this appendix. It consists of human-human interactions for daily motions, such as passing objects, greeting, and communicating, and professional activities, such as, Taekwondo, Latin dance, and boxing. It consists of a total of $7,779$ motions with $22$ joints per person. We randomly select the pose sequence of one of the characters as the acting sequence for each motion to train our model. We aim to synthesize the full body motion of the corresponding other character in each motion as our output. We follow the split of InterGen [42] for our experiments. Since the InterHuman dataset does not have hand motions, we only train with body motions and forego the hand diffusion stage. We train ReMoS for about $45\mathrm{K}$ iterations on the InterHuman dataset using the Adam optimizer [36], with a base learning rate of $10^{-5}$ and a batch size of $64$.
0.C.2 Baseline Preparation
As we mention in Sec. 5.2 in the main paper, we use InterFormer [11], MixNMatch [20], ComMDM [54], RAIG [65] and InterGen [42] as baselines. We describe how we use each of these methods in our setting.
@@ 0.C.2 Baseline Preparation
InterFormer consists of a transformer network with temporal and spatial attentions. It takes an input acting sequence $Y$ and encodes it with spatial and temporal self-attention. It also needs the initial pose of the reactor $X$ and predicts the subsequent frames of the reactor in an autoregressive manner. It uses information from skeletal adjacency matrices and an interaction distance module that provides information on the interactions. We use the normalization technique mentioned in Sec. 5.1 in the main paper to normalize the actor’s and the reactor’s body poses. We train InterFormer on an NVIDIA RTX A4000 GPU for about $20\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $10^{-5}$ and a batch size of $128$. We use $207$ dimensional latent embedding and $6$ layers in the transformer decoder with $3$ heads to calculate the attention.
@@
MixNMatch proposes an end-to-end framework to synthesize stylized reactive motion informed by multi-hot action labels. It operates in one of two settings, interaction mixing and interaction matching. In interaction mixing, it generates a reaction combining different classes of reactive styles according to the multi-label indicator. In interaction matching, it generates the reactive motion corresponding to the interaction type and the input motion. Our setting is similar to interaction matching, where we input the acting sequence into the model and synthesize the reactive sequence. We mask out the action label defining the interaction type from the input and train the reactor’s motion $X$ based on the actor’s motion $Y$. We use the normalization technique mentioned in Sec. 5.1 in the main paper to normalize the actor’s and the reactor’s body poses. We train MixNMatch on an NVIDIA RTX A4000 GPU for about $3.6\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $10^{-5}$ and a batch size of $16$. We use $256$ LSTM neurons for each spatial slice and $1{,}200$ for the attention layer.
@@
Role-Aware Interaction Generation (RAIG) is a diffusion-based model that learns two-person interactions by generating single-person motions for a designated role. The role is supplied in the form of textual annotations, which are translated into active and passive voices to ensure the text is consistent with each role. The model generates interactions with two transformers that share parameters, and a cross-attention module connecting them. The active and passive voice descriptions are proveded as inputs to the corresponding transformers responsible for generating the actor and the reactor. The transformers consist of cross-attention modules both for language and motion. To use RAIG as a baseline for our annotation-free setting, we mask out the cross-attention module for the language in both the transformers and train to generate two-person motions unconditionally. We normalize the interactions as described in the original paper [65]. We train RAIG on an NVIDIA RTX A4000 GPU for about $20\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $2^{-4}$ and a batch size of $32$. We use $512$ dimensional latent embedding and $8$ attention blocks. During inference, we freeze the transformer that learns the actor’s motion $Y$. The other transformer generates the reactor’s motion $X$, being influenced by the actor’s ground truth motion.
@@
InterGen is a diffusion-based approach that generates two-person motions from text prompts. It was originally trained by conditioning on rich textual annotations. It uses cooperative denoisers with novel weight-sharing and a mutual attention mechanism to improve interactions between two persons. To use it as a baseline in our annotation-free setting, we mask out the text embeddings from the model input, and train InterGen to generate two-person motions (both actor and reactor) unconditionally. We use the non-canonical motion representation proposed in the original paper [42]. During inference, we use the customization used in InterGen for person-to-person generation. We take a single-person motion (the actor’s motion $Y$) as input, and freeze it during the forward diffusion process. The frozen weights from the first person propagate into the model, which then uses the ground truth actor’s motions to reconstruct the second person’s motion (the reactor’s motion $X$). We train InterGen on an NVIDIA RTX A4000 GPU for about $30\mathrm{K}$ iterations for both the LindyHop and the Ninjutsu sets of ReMoCap, using the Adam optimizer [36] with a base learning rate of $10^{-4}$, a cosine LR scheduler, and a batch size of $64$.
Appendix 0.D Additional Results
We provide additional results and the trainable parameter counts of all models. We further show how ReMoS can be used as a motion editing tool for character control applications.
0.D.1 Quantitative Evaluation on the InterHuman Dataset [42]
We report additional evaluation of ReMoS compared to its diffusion-based baselines on the InterHuman [42] dataset in Table 0.D.2. We report performance on the standard evaluation metrics, including MPJPE, MPJVE, FID, Diversity and Multi-modality. For Multi-modality, we generate each sequence $5$ times and report numbers with a $95\%$ confidence interval. InterHuman dataset does not provide hand motions, so we only evaluate the reactors’ body motions. ReMoS achieves state-of-the-art performance in the aforementioned metrics in the InterHuman dataset, highlighting the utility of our method for diverse forms of two-person interactions.
Table 0.D.1: Quantitative evaluation on the InterHuman dataset [42]. We compare ReMoS with state-of-the-art motion synthesis methods on the InterHuman [42] dataset. $\downarrow$: lower is better, $\uparrow$: higher is better, $\rightarrow$: values closer to GT are better. Bold indicates best.